OpenStudy (anonymous):

sorry guys I cant get it right please help me im stock ok.. suppose a computer virus begins by infecting 8 computers in the first hour after it released . Each hour after that, each newly infected computer causes 8 more computers to become infected . THE FUNCTION Y=8* MODELS THE SITUATION Make a table with integer values X 1TO 4 PLEASE GUYS THANK YOU

OpenStudy (anonymous):

how many more computers are infected after the first hour?

OpenStudy (anonymous):

THE FUNCTION IS Y=8* MODELS THE SITUATION MAKE A TABLE WITH INTEGER ES OF x 1 to 4

OpenStudy (anonymous):

sorry values of X 1TO 4

OpenStudy (anonymous):

ARE YOU THERE RYAN

OpenStudy (anonymous):

Eight infected computers → eight more infected computers each, to a total of 8*8=64. Then these 64 infect eight more computers each: 64*8=8³. And so on. The function is \(y=8^x\)

OpenStudy (anonymous):

so first line of table is 8^1 which is an easy one = 8

OpenStudy (anonymous):

I DONT UNDERSTAND BU THANK YOU DO ,I NEED MORE SPECIFIC WORDS CUS I NEED TO PUT IN WORDS NOT IN FORMULA

OpenStudy (anonymous):

In the first hour 8 computers are infected: total 8 each of those computers will infect 8 more computers yes? so at time = 2 the total is 64 each of those computer (64) will infect 8 computers 8*64 = 512 at t=3 so the next one would be 512*8 with each tick of the clock each infected computer infects 8 computers which means that we can write that as 8^t with t = the ticks \[8^{t}\] because 8^4 just means 8*8*8*8 = \[8^{4}\]

OpenStudy (anonymous):

WOAO THANK YOU BEEN SO HELPFUL LOL

OpenStudy (anonymous):

THANK YOU HAVE A GOOD NIGHT

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."